OpenStudy (anonymous):

trying to solve lim as x approaches 0 of (h-1)^3+1/h

OpenStudy (whpalmer4):

\[\lim_{h\rightarrow0}\frac{(h-1)^3+1}{h}\] ?

OpenStudy (whpalmer4):

What do you get if you multiply out \((h-1)(h-1)(h-1)\) and then add 1?

OpenStudy (anonymous):

h^3-2h^2+1h^2/-h^3+2h^2-1h^2 =1+1 =2 ?

OpenStudy (whpalmer4):

No, might need to practice that multiplication a bit... \[(h-1)(h-1)(h-1) = (h-1)(h^2-1h-1h+1) =\]\[ (h^3-1h^2-1h^2+1h-1h^2+1h+1h-1)=h^3-3h^2+3h-1\]

OpenStudy (anonymous):

oops-- then if the lim approaches 0 then replacing h with 0 would give me -1

OpenStudy (whpalmer4):

No again. \[\lim_{h\rightarrow0}\frac{(h-1)^3+1}{h} = \lim_{h\rightarrow0}\frac{h^3-3h^2+3h-1+1}{h} \]

OpenStudy (anonymous):

so another try- can i factor out h as in h(h^2-3h+3)/h am i on the right track?

OpenStudy (whpalmer4):

Exactly!

OpenStudy (anonymous):

Thanks!!!!

OpenStudy (whpalmer4):

What was your final answer?

OpenStudy (whpalmer4):

Hopefully, the 3 rd time is the charm ;-)

OpenStudy (anonymous):

3 I think thats it

OpenStudy (whpalmer4):

Yes, that's why I wrote my last message the way I did ;-) 3 is correct.